A normally distributed population has mean 1,214 and standard deviation 122.

a. Find the probability that a single randomly selected element X of the population is between 1,100 and 1,300.

b. Find the mean and standard deviation of 𝑥̅for samples of size 25.

c. Find the probability that the mean of a sample of size 25 drawn from this population is between 1,100 and 1,300.


1
Expert's answer
2021-04-27T00:42:38-0400

We have given that,

"\mu = 1214"

"\sigma = 122"


a.)"P(1100<X<1300)"


"= P(\dfrac{1100-1214}{122}<Z< \dfrac{1300-1214}{122})"


"= P(-0.93<Z<0.70)"


"= 0.70580-0.176"


"= 0.58"


b.) Now we have "n= 25"

Sample mean"= \mu_x = \mu = 1214"

Sample standard devaition "= \sigma_x = \dfrac{\sigma}{\sqrt n} = \dfrac{122}{5} = 24.4"


c.) "P(1100<X<1300)" for "n=25"


"= P(\dfrac{1100-1214}{24.4}<Z< \dfrac{1300-1214}{24.4})"


"= P(-4.67<Z<3.52)"


"= 1"